Hydrazine N2H4 can interact with water in two steps

N2H4(aq) + H2O (I) = N2H5^+(aq) + OH^-(aq) kb1 = 8.5 x 10^-7

N2H5^+(aq) + H2O (I) = N2H6^2+(aq) + OH^-(aq) kb2 = 8.9 x 10^-16

(a) what is the concentration of OH^-, N2H6^2+ in a 0.010 M solution of hydrazine?
(b) what is the pH 0.010 M solution of hydrazine?
1
Expert's answer
2020-11-14T13:53:57-0500

a) Equllibrium concentration"=" "N_2H_4(x)\times N_2H_5^+OH^-(x)"

Equation for equilibrium constant

"=" "[N_2H_6^+][OH^-]\over N_2H_4"

Substituting values in the equation yields;

"x\times x\over0.01-x" "=8.9\times 10^{-16}"

"=x^2+(8.9\times10^{-16}\times0.01M)^{1\over 2}"

Using the quadratic formula;

"a=1
b=8.6\times 10^{-7}
c=-8.9\times 10^{-16}"

"-b \stackrel{+}{-}" "\sqrt {{b^2-4ac}} \over2a"

"=""-8.6\times 10^{-7} \overset{+}{-}" "\sqrt{(8.6\times10^{-7})^2-4(1)(8.9\times 10^{-16}} \over 2(1)"

"=-1.699\times 10^{-6}M"

b) "pH=-log\sqrt{K_b\times C_o}"

"=-log (K_b\times C_o)^{1\over2}"

"=-log (8.9\times10^{-16}\times 0.01M)^{1\over2}"

"=-log 2.983\times 10^{-9}"

"pH=8.5"